OpenStudy (waheguru):

he line (9,-5) is perpinduclar to 3x+y+6=0 what is the equation of the linr (9,-5 he line (9,-5) is perpinduclar to 3x+y+6=0 what is the equation of the linr (9,-5 @Mathematics

OpenStudy (anonymous):

assuming you mean the line through the point (9,-5): slope of the given equation = -3 so slope of rgeuired line is -1/-3 = 1/3 using the general form y-y1 = m(x-x1) y--5 = 1/3 (x-9) y+5 = 1/3 (x-9) 3y+15 = x - 9 3y - x + 24 = 0 is the required equation